# Towards Temporally Consistent Referring Video Object Segmentation

### v1

@@ IV-C3 Temporal Consistency
In Table I, we compare HTR with other open-sourced R-VOS models on Ref-YouTube-VOS using the proposed MCS metric, which evaluates the temporal consistency of segmentation across thresholds from 0.1 to 0.9. HTR significantly outperforms ReferFormer [23], MTTR [24], OnlineRefer [81], and SgMg [22] across all thresholds due to the temporal consistency modeling. When Swin-L is used as the visual encoder, HTR outperforms the baseline ReferFormer by 20.3%, 21.7%, and 52.6% under MCS@0.1, MCS@0.5, MCS@0.9, respectively. Although HTR with VideoSwin-T employs a smaller backbone compared to ReferFormer with Swin-L, it still achieves notably higher MCS. These results prove the effectiveness of HTR in improving temporal consistency. Moreover, the MCS decreases remarkably for thresholds above 0.5, indicating the future research direction of developing R-VOS models that can produce both temporally consistent and accurate segmentation outcomes.
@@ IV-C5 Inference Speed and Parameters
We perform selective conditional segmentation and efficient inter-frame collaboration for temporally consistent R-VOS. By performing conditional segmentation only on top-ranked (reference) frames, i.e., avoiding the slow and computationally heavy cross-modal FPN decoding on the remaining frames, HTR runs 8-16% faster than ReferFormer [23], despite having about 12% (22M) more parameters attributed to the hybrid memory. This demonstrates the great performance trade-off in our approach.
@@ IV-D Qualitative Results
We visualize some typical segmentation results in Fig. 4. ReferFormer and SgMg face challenges when dealing with temporal context changes, such as the target giraffe is not always in the middle, and the existence of other similar-looking objects like giraffes and gorillas. In contrast, HTR predicts correct and temporally consistent object masks, thanks to its temporal consistency modeling. Additionally, SgMg can handle various challenging scenarios including fast motion, occlusion, objects with similar appearance, and small objects, as shown in Fig. 4(b).
@@ IV-E Limitations
Our paradigm relies upon the availability of a few high-quality reference masks. However, the selective segmentation stage may not provide any good reference mask for some challenging videos. One such failure case is shown in Fig. 4(c). In such cases, our paradigm will obviously be unable to improve the segmentation quality. Nevertheless, our approach still significantly outperforms existing methods without using any ground-truth reference mask. Future research could develop more effective selective referring segmentation models.
@@ IV-F Visualization of Global Token Association
TABLE IV: Comparison with well-trained VOS models on Ref-YouTube-VOS. E.: end-to-end. Add. Enc.: additional visual encoder. Add. Data: additional training data. X-HTR: replace our hybrid memory with XMem [10]. HTR is end-to-end and achieves superior performance under fair comparison
@@
IV-G Ablation Study
IV-G1 Well-trained VOS models
In Table IV, we compare HTR with the combination of the benchmark ReferFormer [23] and well-trained temporal propagation models, i.e., CFBI [26], STCN [27], and XMem [10]. Following ReferFormer, we use Swin-L as the visual encoder in this experiment (VideoSwin-T is used in subsequent ablations). We perform comparisons using 360p frames and adopt high-score frames as reference frames for fair comparison. Our end-to-end HTR significantly surpasses others by over 2.7% and runs up to 26$\times$ faster. Moreover, replacing the hybrid memory with XMem in our paradigm results in a 2.1% lower $\mathcal{J}$$&\mathcal{F}$. XMem designs sensory memory to propagate the ground-truth mask of the first frame which does not align with our challenge of propagating imperfect reference masks from any frames. In contrast, HTR extracts robust global context and improves matching and memory extracted from imperfect reference masks. These results confirm the effectiveness of our hybrid memory and paradigm.
TABLE V: Ablation of memory components. Collaborative local and global components achieve the best performance
@@ IV-G1 Well-trained VOS models
TABLE VI: Ablation of memorized ratio
@@
IV-G2 Hybrid Memory
In Table V, we evaluate the individual contributions of the components within our hybrid memory. Both local and global components exhibit strong temporal propagation ability in our paradigm. Notably, global tokens outperform local memory. Our global tokens employ weighted feature aggregation to suppress noise and form robust global representations. This is crucial for improving the temporal consistency and mask quality during propagation especially when the initial reference masks are automatically generated and imperfect. Overall, the collaboration of local and global components leads to superior performance.
IV-G3 Memorized Frames
In Table VI, we evaluate the performance of HTR across different ratios of memorized frames. HTR dynamically selects the top $k$ percent of ranked frames (frame numbers are rounded up) to construct the hybrid memory. The results show that memorizing the top 25% of ranked frames to segment the remaining 75% achieves the best performance.
TABLE VII: Ablation of strides
@@ IV-G4 Different Feature Strides
In this work, HTR performs inter-frame collaboration using features with stride 16, a common setting for memory modules [64, 10]. Table VII presents our evaluation of different strides, demonstrating that the common setting achieves the best performance.
TABLE VIII: Integrate HTR to large R-VOS model
[…]
UNINEXT [79] 70.1 67.5 72.6
Ours 72.5 70.2 74.8
Figure 6: Our model (GTR. in the figure) is the top performer on the active 4th R-VOS competition leaderboard, captured in May 2023.
IV-H Collaboration with Large R-VOS Model
Our trained HTR models can seamlessly collaborate with other large R-VOS models to boost temporal consistency and performance even without re-training since inter-frame collaboration takes general masks as input. We leverage UNIEXT [79] to generate reference masks for our trained model. We refer to this variant as UNI-HTR for brevity. UNIEXT is a recent large model that achieves impressive performance. It uses the ViT-Huge [105] visual encoder and 32 A100 GPUs to train on 11 large-scale training datasets from related tasks. In Table VIII, our approach improves the performance by 2.4% on Ref-YouTube-VOS without techniques like multi-scale inference and model ensembles. This leads to top position on the active leaderboard in the 4th R-VOS competition on Codalab, as evidenced in Fig. 6 (captured in May 2023).

### v2

@@ IV-C3 Temporal Consistency
In Table I, we compare HTR with other open-sourced R-VOS models on Ref-YouTube-VOS using the proposed MCS metric, which evaluates the temporal consistency of segmentation across thresholds from 0.1 to 0.9. HTR significantly outperforms ReferFormer [22], MTTR [23], OnlineRefer [80], SgMg [21], and SOC [98] across all thresholds due to the temporal consistency modeling. When Swin-L is used as the visual encoder, HTR outperforms the baseline ReferFormer by 20.3%, 21.7%, and 52.6% under MCS@0.1, MCS@0.5, MCS@0.9, respectively. Although HTR with VideoSwin-T employs a smaller backbone compared to ReferFormer with Swin-L, it still achieves notably higher MCS. These results prove the effectiveness of HTR in improving temporal consistency. Moreover, the MCS decreases remarkably for thresholds above 0.5, indicating the future research direction of developing R-VOS models that can produce both temporally consistent and accurate segmentation outcomes.
@@ IV-C5 Inference Speed and Parameters
We perform selective conditional segmentation and efficient inter-frame collaboration for temporally consistent R-VOS. By conducting conditional segmentation only on top-ranked (reference) frames, i.e., avoiding intensive cross-modal FPN decoding on remaining frames, HTR with VideoSwin-T achieves 58 FPS on a single RTX 3090 GPU. This is faster than ReferFormer [22] (50 FPS) and comparable to the recent SOC [98] (59 FPS) using the same backbone. Additionally, HTR with VideoSwin-B and Swin-L runs at 36 and 42 FPS, respectively. Despite having approximately 12% (22M) more parameters due to the hybrid memory, our model runs faster and outperforms ReferFormer by up to 5% points on Ref-YouTube-VOS, demonstrating good performance trade-off.
TABLE IV: Quantitative comparison on MeViS
[…]
URVOS [35] 27.8 25.7 29.9
LBDT [75] 29.3 27.8 30.8
MTTR [23] 30.0 28.8 31.2
ReferFormer [22] 31.0 29.8 32.2
VLT+TC [100] 35.5 33.6 37.3
LMPM [100] 37.2 34.2 40.2
HTR (Ours) 42.7 39.9 45.5
IV-C6 Results on MeViS
In Table IV, we further evaluate HTR with the tiny version backbone on the recent MeViS [100] dataset. Training from scratch on the MeViS train split, our model achieves 42.7 $\mathcal{J}$$&\mathcal{F}$, outperforming ReferFormer by 11.7% points and demonstrating advanced performance. This improvement is more significant because the MeViS dataset focuses on moving objects and longer videos, challenging previous models with temporal consistency issues.
@@ IV-D Qualitative Results
We visualize some typical segmentation results in Fig. 4. ReferFormer and SgMg face challenges when dealing with temporal context changes, such as the target giraffe is not always in the middle, and the existence of other similar-looking objects like giraffes and gorillas. In contrast, HTR predicts correct and temporally consistent object masks, thanks to its temporal consistency modeling. Additionally, HTR can handle appearance and motion expressions in various challenging scenarios, including fast motion, occlusion, objects with similar appearance, and small objects.
@@ IV-E Limitations
Our paradigm relies upon the availability of a few high-quality reference masks. However, the selective segmentation stage may not provide any good reference mask for some challenging videos. Two such failure cases are shown in Fig. 4(c). In such cases, our paradigm will obviously be unable to improve the segmentation quality. Nevertheless, our approach still significantly outperforms existing methods without using any ground-truth reference mask. Future research could develop more effective selective referring segmentation models.
@@ IV-F Visualization of Global Token Association
TABLE V: Comparison with well-trained VOS models on Ref-YouTube-VOS. E.: end-to-end. Add. Enc.: additional visual encoder. Add. Data: additional training data. X-HTR: replace our hybrid memory with XMem [10]. HTR is end-to-end and achieves superior performance under fair comparison
@@
TABLE VI: Ablation of memory components. Collaborative local and global components achieve the best performance
@@
TABLE VII: Ablation of memorized ratio
@@
TABLE VIII: Ablation of strides
@@
TABLE IX: Ablation of number of training frames
[…]
3 51.0 62.0 60.6 63.4
5 53.8 64.2 62.7 65.7
10 54.6 64.9 63.4 66.4
TABLE X: Performance on videos with different numbers of frames
Method $0<N\leq 50$ $50<N\leq 80$ $N>80$
ReferFormer 59.8 59.4 61.5
HTR (Ours) 66.9 63.1 67.0
IV-G Ablation Study
IV-G1 Well-trained VOS models
In Table V, we compare HTR with the combination of the benchmark ReferFormer [22] and well-trained temporal propagation models, i.e., CFBI [25], STCN [26], and XMem [10]. Following ReferFormer, we use Swin-L as the visual encoder in this experiment (VideoSwin-T is used in subsequent ablations). We perform comparisons using 360p frames and adopt high-score frames as reference frames for fair comparison. Our end-to-end HTR significantly surpasses others by over 2.7% and runs up to 26$\times$ faster. Moreover, replacing the hybrid memory with XMem in our paradigm results in a 2.1% lower $\mathcal{J}$$&\mathcal{F}$. XMem designs sensory memory to propagate the ground-truth mask of the first frame which does not align with our challenge of propagating imperfect reference masks from any frames. In contrast, HTR extracts robust global context and improves matching and memory extracted from imperfect reference masks. These results confirm the effectiveness of our hybrid memory and paradigm.
IV-G2 Hybrid Memory
In Table VI, we evaluate the individual contributions of the components within our hybrid memory. Both local and global components exhibit strong temporal propagation ability in our paradigm. Notably, global tokens outperform local memory. Our global tokens employ weighted feature aggregation to suppress noise and form robust global representations. This is crucial for improving the temporal consistency and mask quality during propagation especially when the initial reference masks are automatically generated and imperfect. Overall, the collaboration of local and global components leads to superior performance.
IV-G3 Memorized Frames
In Table IX, we evaluate the performance of HTR across different ratios of memorized frames. HTR dynamically selects the top $k$ percent of ranked frames (frame numbers are rounded up) to construct the hybrid memory. The results show that memorizing the top 25% of ranked frames to segment the remaining 75% achieves the best performance.
@@ IV-G4 Different Feature Strides
In this work, HTR performs inter-frame collaboration using features with stride 16, a common setting for memory modules [63, 10]. Table IX presents our evaluation of different strides, demonstrating that the common setting achieves the best performance.
IV-G5 Number of Training Frames
We follow previous works such as [22, 21, 79] to randomly sample 5 frames during training for fair comparison. This random sampling enables the model to learn both long-term and short-term matching for temporally consistent R-VOS. In Table IX, we also explore different numbers of training frames. Our model consistently performs well, with $N$=5 demonstrating the best trade-off.
IV-G6 Videos with Different Numbers of Frames
In Table X, we compare the performance ($\mathcal{J}$$&\mathcal{F}$) of our model and the baseline ReferFormer on Ref-DAVIS17 videos with varying frame numbers. HTR achieves significantly better results than ReferFormer across both short and long videos.